OpenStudy (anonymous):

Solve this differential equation by eliminating the arbitrary constants. y=Acos(log x) + Bsin(log x)

OpenStudy (anonymous):

the equation or differential equation?

OpenStudy (anonymous):

diffrential eqn

OpenStudy (anonymous):

but where are the differentials?

OpenStudy (anonymous):

you have to eliminate A & B

OpenStudy (anonymous):

i guess this is the prob to form a DE

OpenStudy (anonymous):

Yea DE.

OpenStudy (anonymous):

question makes no sense

OpenStudy (anonymous):

eq contains 2 constants, differentiate twice n solve

OpenStudy (anonymous):

how this makes no sense? this is que from book. @ elecengineer

OpenStudy (anonymous):

Yea But M unable to solve.can u help me in differentiating?

OpenStudy (anonymous):

y'=-Asin(logx)(1/x)+Bcos(logx)(1/x) y''=-Acos(logx)(-1/x^2)-Bsin(logx)(-1/x^2) =(1/x^2)[Acos(logx)+Bsin(logx)] y''=(1/x^2)y

OpenStudy (anonymous):

okay thank u so much uzma let me check

OpenStudy (anonymous):

what :|